# Position-Aware Depth Decay Decoding (): Boosting Large Language Model Inference Efficiency

### v1

@@ 4.3 Factor Study
Firstly, we examine the effect of decay rate and Flex Layer depth on performance on GSM8K, as shown in Figure 6. We draw the following conclusions: (i) At the same decay rate, it is more appropriate to start the decay from the middle of the model for models of various sizes. This experimental conclusion confirms our hypothesis in Section 3.3 that LLMs have Core and Flex Layers. The Core Layer generally represents the layers at the beginning and end, while the Flex Layer represents the middle layers, where redundant layers exist and can be skipped to improve inference efficiency. (ii) Additionally, the results presented in the technical appendix for BBH exhibit a similar trend, with optimal hyperparameters differing from those for GSM8K. This observation suggests that and are task-specific parameters. We hypothesize that this is due to the BBH task having a much shorter output length compared to GSM8K, which necessitates a faster decay rate (i.e., a smaller decay coefficient) for optimal performance.
@@ 6 Limitation
Appendix A Specific Values
In this appendix, we provide the detailed numerical results for the BBH benchmark, as shown in Table 8. Table 9 presents the description and search range for each parameter during the hyperparameter tuning process. Table 9 demonstrates the successful transfer of optimal parameters, identified using 10% of the validation set, from smaller models to larger-scale models.

### v2

@@ 4.3 Factor Study
Firstly, we examine the effect of decay rate and Flex Layer depth on performance on GSM8K, as shown in Figure 6. We draw the following conclusions: (i) At the same decay rate, it is more appropriate to start the decay from the middle of the model for models of various sizes. This experimental conclusion confirms our hypothesis in Section 3.3 that LLMs have Core and Flex Layers. The Core Layer generally represents the layers at the beginning and end, while the Flex Layer represents the middle layers, where redundant layers exist and can be skipped to improve inference efficiency. (ii) Additionally, the results presented in the Appendix Table 9 for BBH exhibit a similar trend, with optimal hyperparameters differing from those for GSM8K. In this appendix, we provide the detailed numerical results for the BBH benchmark, as shown in Table 8. Table 9 presents the description and search range for each parameter during the hyperparameter tuning process. Table 10 demonstrates the successful transfer of optimal parameters, identified using 10% of the validation set, from smaller models to larger-scale models.
This observation suggests that and are task-specific parameters. We hypothesize that this is due to the BBH task having a much shorter output length compared to GSM8K, which necessitates a faster decay rate (i.e., a smaller decay coefficient) for optimal performance.
@@ 6 Limitation
Acknowledgments
This work is supported by the National Science and Technology Major Project (No. 2022ZD0116314), the National Science Foundation of China (No. 62106249).